OpenStudy (anonymous):

Two cards are drawn in succession without replacement from a standard deck of 52 cards. What is the probability that the first card is a spade given that the second card is a diamond? (Round your answer to three decimal places.)

OpenStudy (anonymous):

I have an idea on how to do this but i think im messsing up somewhere along the process

OpenStudy (anonymous):

are you familiar with bayes theorem?

OpenStudy (anonymous):

yes.. that is what I am using, but I cant seem to get the right answer. I dont know what I am doing wrong

OpenStudy (anonymous):

alright lets see what we can do.. D = 2nd card is diamond S = 1st card is spade \[P(S | D) = \frac{ P(D|S)P(S) }{ P(D) }\]

OpenStudy (anonymous):

so here i think the P(D | S) = (13/52)(13/52) and P(S) = 13/52 but im not quite sure on how to get the P(D)

OpenStudy (anonymous):

sorry p(D|S) = (13/52)(13/51)

OpenStudy (anonymous):

well i used the other form of baye theorem \[P(S|D) = \frac{ P(D|S)P(S) }{ P(D|S)P(S) + P(D|S')P(S') }\] where \[P(D|S') = {(\frac{ 52-13 }{ 52 })\frac{ 13 }{ 51 } }\] and P(S') = 39/52

OpenStudy (anonymous):

ya i was using P(D) = (39/52)(13/51) + (13/52)(13/51)

OpenStudy (anonymous):

but it came out as a low number which it feels kinda wrong unless its correct lol

OpenStudy (anonymous):

what did you get?

OpenStudy (anonymous):

i was getting .250 and it was wrong

OpenStudy (anonymous):

i got around 10% assuming did the algebra correctly, i did it fast so idk if i made a mistake or not

OpenStudy (anonymous):

let me calculate it again

OpenStudy (anonymous):

I got 0.1 = 10% its pretty low so it feels kinda wrong lol

OpenStudy (anonymous):

it has to be rounded to 3 decimal places... do you know the next 2 numbers?

OpenStudy (anonymous):

its just 0.1, so 0.100

OpenStudy (anonymous):

no its not right :(

OpenStudy (anonymous):

yea thought so i didnt think it was right but idk what exactly im doing wrong

OpenStudy (anonymous):

try 0.255 see if it's right

OpenStudy (anonymous):

yes!!! that problem was driving me crazy!

OpenStudy (anonymous):

thanks!